Rubbing alcohol contains 585 grams of isopropanol (C3H7OH) per liter of solution. Calculate the molarity.
1
Expert's answer
2020-01-30T05:01:36-0500

mass =585g

molar mass of C3H7OH= 60.1

number of moles= 585/60.1 =9.734moles

"molarity=moles\/litre"

molarity= 9.734/1000 =0.0097M